# Interactive Visualization Recommendation with Hier-SUCB

### v2

@@ Abstract.
journalyear: 2025copyright: rightsretainedconference: Proceedings of the ACM Web Conference 2025; April 28–May 2, 2025; Sydney, NSW, Australia.booktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28–May 2, 2025, Sydney, NSW, Australiaisbn: 979-8-4007-1274-6/25/04doi: 10.1145/3696410.3714697ccs: Information systems Personalizationccs: Human-centered computing Visualization systems and tools
@@ 4. Algorithm & Theoretical Analysis
In this section, we propose our combinatorial bandit algorithm for interactive personalized visualization recommendation, called Hierarchical Semi UCB (Hier-SUCB). The code is available at https://github.com/HierSUCB-WWW/HierSUCB/tree/main.
@@ 5.1. Experimental Setup
In this section, we discuss our experimental setup. First, we introduce the baseline algorithms, then discuss the metrics used for evaluation.
@@ 5.1.2. Metric
We follow C2UCB  (Qin et al., 2014) in the definition of cumulative regret and average reward. We get average reward of one round by computing the cumulative sum of the mean of the user feedback in each round. Similarly, cumulative regret is computed from the cumulative sum of the mean regret in each round. When comparing to the offline method Neuro-PVR  (Qian et al., 2022), we compare the HR@K over iterations. Recall that HR@K (hit rate at $k$) is the fraction of the top $k$ recommended visualizations that are in the set of visualizations that are actually relevant to the user. We compare our method with the offline method Neuro-PVR  (Qian et al., 2022) as well as with other bandit algorithms. For this comparison, we use HR@1 since the bandit algorithms mentioned are designed for recommending one visualization. Furthermore, HR@1 is naturally the most important, since it indicates how likely the approaches are to recommend the desired visualization to the user directly without showing a sequence of visualizations.
@@ 5.2.4. User study of simulator
To validate the performance of simulator in the real life, we conducted a user study based on the visualization generated by algorithms. We first recorded the visualization generated by Hier-SUCB and C2UCB given the feedback of the simulator. Then we recruited 51 participants (university students aged from 18 to 26, 47 of them reported experience of creating visualizations) and recorded their preference on 10 samples randomly selected from those visualizations. As shown in Fig.  3 (a), we compared the percentage of user preference in round 20 and round 50 of the visualization generated in simulated experiments. From the human evaluation, we observe an increase in the population of participants that like either of the visualization, which implies the simulator can help bandit algorithms learn user preferences in the real-world setting.

### v5

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714697isbn: 979-8-4007-1274-6/2025/04ccs: Information systems Personalizationccs: Human-centered computing Visualization systems and tools
@@ 4. Algorithm & Theoretical Analysis
In this section, we propose our combinatorial bandit algorithm for interactive personalized visualization recommendation, called Hierarchical Semi UCB (Hier-SUCB).
@@ 5.1. Experimental Setup
In this section, we discuss our experimental setup. First, we introduce the baseline algorithms, and then metrics used for evaluation.
@@ 5.1.2. Metric
We follow C2UCB  (Qin et al., 2014) in the definition of cumulative regret and average reward. We get average reward of one round by computing the cumulative sum of the mean of the user feedback in each round. Similarly, cumulative regret is computed from the cumulative sum of the mean regret in each round. When comparing to the offline method Neuro-PVR  (Qian et al., 2022), we compare the HR@K over iterations. Recall that HR@K (hit rate at $k$) is the fraction of the top $k$ recommended visualizations that are in the set of visualizations that are actually relevant to the user. We compare our method with the offline method Neuro-PVR  (Qian et al., 2022) as well as with other bandit algorithms. For comparison, we use HR@1 as the bandit algorithms mentioned are designed for recommending one visualization. Furthermore, HR@1 is naturally the most important, since it indicates how likely the approaches are to recommend the desired visualization to the user directly.
@@ 5.2.4. User study of simulator
To validate the performance of simulator in the real life, we conducted a user study based on the visualization generated by algorithms. We first recorded the visualization generated by Hier-SUCB and C2UCB given the feedback of the simulator. Then we recruited 51 participants (university students aged from 18 to 26, 47 of them reported experience of creating visualizations) and recorded their preference on 10 samples randomly selected from those visualizations. As shown in Fig.  3 (a), we compared the percentage of user preference in round 20 and 50 of the visualization generated in simulated experiments. From human evaluation, we observed an increase in the percentage that like either of the visualization, which implies the simulator can help bandit algorithms learn user preferences in the real-world setting.